# Scalable Sensor Placement for Cyclic Networks with Observability Guarantees: Application to Water Distribution Networks

### v1

@@ IV Implementation on EPANET benchmarks
For the sensor placement Algorithms 1 and 2, a structured representation of a Water Distribution Network (WDN) model is required. To this end, we utilize the linearized elastic water column (EWC) model, as formalized in [25] and given by
[…]
To perform sensor placement defined in Section III, we utilize the structured system representation of this model. This structured form, as formalized in [20, Section IV.A], is given by
$\mathcal{A}_{\mathrm{WDN}}=\begin{bmatrix}\mathcal{I}&\mathcal{A}_{\mathrm{inc}}^{\top}\\ \mathcal{A}_{\mathrm{inc}}&\mathcal{D}\\ \end{bmatrix}\!\!$, (8)
with the components defined as
$\mathcal{I}$ $=\text{diag}\{\ast\}\in\{0,\ast\}^{m\times m},$
$\mathcal{A}_{\mathrm{inc}}(i,j)$ $=\begin{cases}0&\text{if }A(i,j)=0\\ \ast&\text{otherwise,}\end{cases}$
$\mathcal{D}$ $=\text{diag}\{?\}\in\{0,?\}^{n\times n}.$
It is important to note that for most linearized WDN models, the off-diagonal blocks in (8) remain unchanged due to the underlying diffusive coupling between pressure and flow. Specifically, the pressure drop across a pipe between nodes $i$ and $j$ follows
$h_{i}-h_{j}=R_{ij}|q_{ij}|^{\alpha-1}q_{ij},$
where $R_{ij}$ is the pipe resistance and $\alpha$ is a model-dependent exponent, e.g., $\alpha=2$ for Darcy-Weisbach or $\alpha=1.85$ for Hazen-Williams [25, 26, 27]. The linearization around an operating point alters only the numerical values, not the structure. As a result, the off-diagonal structure in (8), which defines the network connectivity, remains the same. Structural variations may still occur in the diagonal entries, reflecting physical properties such as resistance, valves, or other hydraulic elements [25], but these do not affect the sensor placement introduced in Section III. Therefore, the proposed method yields the same sensor configuration for any WDN model that incorporates this form of coupling.
@@
To implement the proposed sensor placement in MATLAB, particularly for EPANET networks, we first construct the incidence matrix $A_{\mathrm{inc}}$, which is required for the structured WDN model (8). Algorithm 3 utilizes the EPANET network link data (lines 3-4) to generate the incidence matrix $A_{\mathrm{inc}}$ (lines 6-10). Using $A_{\mathrm{inc}}$, we define the structured system dynamics (8), and implement the system in MATLAB using symbolic variables star and Question. The computational complexity of Algorithm 3 is $\text{O}(n_{x})$, determined by the for loop in lines 6–10.
To illustrate the structured modeling approach and its associated graph, consider the example WDN shown in Fig.5. From the corresponding incidence matrix
$A_{\mathrm{inc}}\!=\!\begin{bmatrix}-1\!&\!1\!&\!1\!&\!0\\ 0\!&\!0\!&\!-1\!&\!1\\ 0\!&\!-1\!&\!0\!&\!-1\\ 1\!&\!0\!&\!0\!&\!0\\ \end{bmatrix}\!$,
we construct the structured matrix in (8). The resulting graph $\mathcal{G}(\mathcal{A}_{\mathrm{WDN}}^{\top})$ is shown in Fig.6.
As can be seen in Figs.5 and6, this graph reflects the physical topology but represents each state variable, flow or pressure, as a separate node. In this example, nodes $q_{1}$ to $q_{4}$ represent flows through pipes $e_{1}$ to $e_{4}$, while nodes $h_{1}$ to $h_{4}$ correspond to pressures at junctions 1 to 3 and tank 4. Self-loops on flow nodes represent head loss due to friction, while self-loops on pressure nodes capture potential pressure losses from valves or other local hydraulic effects [25].
Figure 5: Topology of a triangular example WDN.
Figure 6: Graph representation $\mathcal{G}(\mathcal{A}_{\mathrm{WDN}}^{\top})$ derived from the structured WDN model corresponding to the example water network in Fig. 5.
Having illustrated the structured modeling approach and its graph representation on a small example, we now demonstrate the scalability and effectiveness of the proposed sensor placement method on larger benchmark networks. The results of applying the spanning tree Algorithm 1 in combination with the sensor placement algorithm are shown in Table I [^1] All experiments were conducted on a Lenovo ThinkPad with an Intel Core i7-165H vPro, 64GB RAM, running MATLAB R2023a..
@@
As can be seen in Table I, the method computes sensor configurations in under 0.1 seconds, even for the largest networks, demonstrating strong computational efficiency. The number of sensors required depends mostly on the number of extreme nodes, the number of cycles, and the network’s connectivity. In most cases, e.g., Hanoi, Net3, and D-town, sensors are placed at extreme nodes, with one additional sensor per cycle, resulting in a total count equal to the sum of extreme nodes and cycles. This aligns with the structure of the sensor placement algorithm, which disconnects cycles and places sensors at the resulting extreme nodes, thereby ensuring at least one sensor per cycle. However, in more interconnected networks with overlapping cycles and fewer extreme nodes, such as AnyTown and L-Town, additional sensors may be needed within cycles.
@@
In Figures 7 and 9, the Hanoi and D-town networks are visualized while Figures 8 and 10 show the corresponding sensor placements using Algorithms 1 and 2. For the Hanoi network, sensor nodes are placed at extreme nodes and near intersection nodes, consistent with the proof of Theorem III.5. The D-town network illustrates the complexity of large-scale cyclic structures and emphasizes the need for scalable sensor placement strategies.

### v2

@@ IV Implementation on EPANET benchmarks
For the sensor placement Algorithms 1 and 2, a structured representation of a Water Distribution Network (WDN) model is required. Therefore, for all examples, we utilize the linear elastic water column (EWC) model, as described in [25], with the associated structured system outlined in [20, Section IV.A] to capture the dynamics of the WDNs. To implement this in MATLAB, particularly for EPANET networks, we define a function to construct the incidence matrix $A_{\mathrm{inc}}$, which is needed for the structured WDN model. Algorithm 3 utilizes the EPANET network link data (lines 3-4) to generate the incidence matrix $A_{\mathrm{inc}}$ (lines 6-10). Using $A_{\mathrm{inc}}$, we define the structured system dynamics as detailed in [20, Section IV.A], and implement the system in MATLAB using symbolic variables star and Question. The computational complexity of Algorithm 3 is $\text{O}(n_{x})$, determined by the for loop in lines 6–10.
@@
The results of the implemented spanning tree Algorithm 1 combined with the sensor placement Algorithm are shown in Table I [^1] All experiments were conducted on a Lenovo ThinkPad with an Intel Core i7-165H vPro, 64GB RAM, running MATLAB R2023a.. As can be seen in Table I, the proposed algorithms compute a sensor placement configuration in under 0.1 seconds, even for the largest network, demonstrating its scalability. The number of placed sensors depends on the number of extreme nodes, cycles, and network topology. In most cases, e.g., Hanoi, Net3, and D-town, sensors are placed at extreme nodes, with one additional sensor per cycle, resulting in a total count equal to the sum of extreme nodes and cycles. This aligns with the structure of the sensor placement algorithm, which disconnects cycles and places sensors at the resulting extreme nodes, thereby ensuring at least one sensor per cycle. However, in more interconnected networks with overlapping cycles and fewer extreme nodes, such as AnyTown and L-Town, additional sensors may be needed within cycles.
@@
In Figures 5 and 7, the Hanoi and D-town networks are visualized, while Figures 6 and 8 show the corresponding sensor placements using Algorithms 1 and 2. For the Hanoi network, sensor nodes are placed at extreme nodes and near intersection nodes, consistent with the proof of Theorem III.5. The D-town network illustrates the complexity of large-scale cyclic structures and emphasizes the need for scalable sensor placement strategies.